# Embedded structure in quantum theory, functional operator and multiverse

### v2

@@ Abstract
We explore a wider theoretical framework that has quantum field theory built-in, taking the fact that quantum mechanics is reconstructed from quantum field theory as a hint. We formulate a quantum theory with an embedded structure by introducing functional operators, and we find that it could describe the level II multiverse. Topics related to a beginning of the universe such as inflation, the third quantization and the landscape are discussed in our formulation.
@@ 1 Introduction
Quantum field theory (QFT) or quantum theory of fields offers an excellent framework that explains a great variety of phenomena in our world and describes curious features such as creation and annihilation of particles, the dichotomy between a particle and a wave, and so on [1, 2]. In fact, QFT has been applied to various systems and been massively successful. For example, particle physics at the electroweak scale is excellently explained by the standard model [3], and the superconductivity becomes better understood by BCS theory [4] in condensed matter physics.
@@
    What is an origin of particles and spacetime? Which came first particles or space-time?
@@
In this paper, we explore a wider theoretical framework that has QFT built-in, taking the fact that QM is reconstructed from QFT as a hint. We formulate a quantum theory with an embedded structure by introducing functional operators, and we find that it could describe the level II multiverse. Topics related to a beginning of the universe such as inflation, the third quantization and the landscape are discussed in our formulation.
@@ 2.1 Framework of quantum field theory
where $\widehat{\varphi}(\bm{x},t)$ is a quantum field, $\bm{x}=(x^{1},x^{2},x^{3})$ and $t$ stand for coordinates of space and time, respectively, $\hbar$ is reduced Planck constant, and $\widehat{H}$ is the Hamiltonian operator in QM containing $\bm{x}$ and its derivatives $\bm{\nabla}$, i.e., $\widehat{H}=\widehat{H}(\bm{x},-i\hbar\bm{\nabla})$.
For example, for an electron with a mass $m_{\rm e}$ in non-relativistic QM, $\widehat{H}$ is given by
$\widehat{H}=-\frac{\hbar^{2}}{2m_{\rm e}}\bm{\nabla}^{2}+V(\bm{x}),$ (2)
@@
where $c$ is a speed of light, $\bm{\alpha}=(\alpha^{1},\alpha^{2},\alpha^{3})$ and $\beta$ are $4\times 4$ Hermitian matrices satisfying $\alpha^{i}\alpha^{j}+\alpha^{j}\alpha^{i}=2\delta^{ij}I$ ($i,j=1,2,3$,$I$:4×$\times 4$ 4 unit matrix), $\alpha^{i}\beta+\beta\alpha^{i}=0$ and $\beta^{2}=I$.
@@
We impose the following quantization conditions on field operators,
@@ 2.2 Derivation of quantum mechanics
using eq. (14), $\langle 0|e^{\frac{i}{\hbar}\widehat{H}_{\varphi}t}=\langle 0|$, i.e., $\langle 0|\widehat{H}_{\varphi}=0$ and eq. (10). From eq. (24) and the first equation in eq. (8), we find that $\psi(\bm{x},t)$ satisfies the Schrödinger equation in QM:
@@
from the rotation. In this way, it is ascertained that a nested structure exists for $\widehat{P}_{\mu}$ and $\widehat{\bm{J}}$ other than the Hamiltonian operator, as seen from the formulas (33) and (34).
@@
we arrive at a probability interpretation that $\widetilde{\psi}^{\dagger}(\bm{k},t)\widetilde{\psi}(\bm{k},t)$ is the probability density and $\langle\psi(t)|\widehat{\bm{p}}|\psi(t)\rangle$ is the expectation value of momenta for the one-particle system in QM.
@@
In the Schödinger picture, observables in QFT are, in general, given as a form
@@
with a potential $\widehat{V}_{\rm I}(t)$, using the Born approximation in a non-relativistic limit.
We list central features concerning the relationship between QFT and QM.
@@ 3 Quantum field’s functional theory
Let us come back to Q1, i.e., “Why does QFT work as an effective theory of elementary particles extremely well?  Why are particles or fields quantized in the first place?”. In concrete, why is a fermion described by the field operator such as $\widehat{\pi}(\bm{x})=i\hbar{\delta}/{\delta{\varphi}(\bm{x})}$, i.e., $\widehat{\varphi}^{\dagger}(\bm{x})={\delta}/{\delta{\varphi}(\bm{x})}$, using the functional derivative?
If there were a fundamental framework that reach QFT in the similar way as the derivation of QM from QFT, we can have an answer to the above questions. To explore such a framework, we will draw on the embedded structure in QFT, e.g.,
@@ 3.1 Framework of quantum field’s functional theory
where $\widehat{\varPhi}(\{\varphi\},t)$ is a functional operator, $\{\varphi\}$ and $t$ stand for a field of $\varphi$ and time, respectively, and $\widehat{H}_{\varphi}$ is the Hamiltonian operator in QFT containing $\varphi(\bm{x})$ and its functional derivatives $\delta/\delta\varphi(\bm{x})$ (see eq. (16)). For instance, $\widehat{H}_{\varphi}$ is given by
@@
We impose the following quantization conditions on functional operators333 If eigenvalues of $\widehat{H}_{\varphi}$ are bounded below, this system can be quantized using commutation relations although the Lagrangian consists of a term with a first-order time derivative and a (sign-flipping of) Hamiltonian (see eq. (52)). This feature is different from that of the quantization for particles with spin 1/2 in QFT. In fact, when we impose commutation relations on Dirac or Weyl fermion, the system becomes ill-defined with the advent of negative norm states. It stems from the existence of negative-energy states in relativistic QM.,
@@ 3.2 Derivation of quantum field theory
  • Elementary particles create and annihilate. Particles can appear in the vacuum, and a multi-particle state can return to the vacuum state. A vacuum state plays a role as a basis to construct any physical states, e.g., multi-particle states are obtained by multiplying creation operators.
@@
${\varPsi}(\{\varphi\},t)=\psi^{(0)}(t)+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}~{}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N}),$ (75)
@@
$\widehat{\varPhi}(\{\varphi\},t)=\widehat{\varPhi}(\{\varphi\})e^{-\frac{i}{\hbar}\mathscr{E}_{0}t},~{}~{}\widehat{\varPhi}^{\dagger}(\{\varphi\},t)=\widehat{\varPhi}^{\dagger}(\{\varphi\})e^{\frac{i}{\hbar}\mathscr{E}_{0}t}.$ (79)
@@
Next let us justify that a fermion field $\varphi(\bm{x})$ is quantized to satisfy the anti-commutation relations (12). Under an infinitesimal translation $\bm{x}^{\prime}=\bm{x}-\bm{\varepsilon}$, the field $\varphi(\bm{x})$ transforms as $\varphi^{\prime}(\bm{x}^{\prime})=\varphi^{\prime}(\bm{x}-\bm{\varepsilon})=\varphi(\bm{x})$, and then an infinitesimal change of $\varphi(\bm{x})$ is given by
@@
In a same way, the expectation value of energy for the state given by eq. (74) is calculated as
${}^{(1)}\langle\varPsi_{\rm M}(t)|\widehat{H}_{\varPhi}|\varPsi_{\rm M}(t)\rangle^{(1)}=\langle\varPsi(t)|\widehat{H}_{\varPhi}|\varPsi(t)\rangle$
$~{}~{}~{}~{}={\varPsi}^{\dagger}(\{\varphi\},t)\widehat{H}_{\varphi}{\varPsi}(\{\varphi\},t)=\int\mathscr{D}\varphi~{}{\varPsi}^{\dagger}(\varphi,t)\widehat{H}_{\varphi}{\varPsi}(\varphi,t)$
$~{}~{}~{}~{}=\langle\varPsi(t)|\widehat{H}_{\varphi}|\varPsi(t)\rangle.$ (89)

### v3

@@ Abstract
We explore a wider theoretical framework that has quantum field theory built-in, taking the fact that quantum mechanics is reconstructed from quantum field theory as a hint. We formulate a quantum theory with an embedded structure by introducing functional operators, and we find that it could describe the level II multiverse. Topics related to a beginning of the universe such as an inflation, the third quantization and the landscape are discussed in our formulation.
@@ 1 Introduction
Quantum field theory (QFT) or quantum theory of fields offers an excellent framework that explains a great variety of phenomena in our world and describes curious features such as a creation and annihilation of particles, the dichotomy between a particle and a wave, and so on [1, 2]. In fact, QFT has been applied to various systems and been massively successful. For example, particle physics at the electroweak scale is excellently explained by the standard model [3], and the superconductivity becomes better understood by BCS theory [4] in condensed matter physics.
@@
    What is an origin of particles (fields) and spacetime? Which came first particles or space-time?
@@
In this paper, we explore a wider theoretical framework that has QFT built-in, taking the fact that QM is reconstructed from QFT as a hint. We formulate a quantum theory with an embedded structure by introducing functional operators, and we find that it could describe the level II multiverse. Topics related to a beginning of the universe such as an inflation, the third quantization and the landscape are discussed in our formulation.
@@ 2.1 Framework of quantum field theory
where $\widehat{\varphi}(\bm{x},t)$ is a quantum field, $\bm{x}=(x^{1},x^{2},x^{3})$ and $t$ stand for coordinates of space and time, respectively, $\hbar$ is the reduced Planck constant, and $\widehat{H}$ is the Hamiltonian operator in QM containing $\bm{x}$ and its derivatives $\bm{\nabla}$, i.e., $\widehat{H}=\widehat{H}(\bm{x},-i\hbar\bm{\nabla})$. Here, we choose the Lagrangian density with a first time-derivative term, because the compatibility (relationship) with QM is easily comprehensible, as will be seen.
For example, for a particle with a mass $m$ in non-relativistic QM, $\widehat{H}$ is given by
$\widehat{H}=-\frac{\hbar^{2}}{2m}\bm{\nabla}^{2}+V(\bm{x}),$ (2)
@@
where $c$ is the speed of light, $\bm{\alpha}=(\alpha^{1},\alpha^{2},\alpha^{3})$ and $\beta$ are $4\times 4$ Hermitian matrices satisfying $\alpha^{i}\alpha^{j}+\alpha^{j}\alpha^{i}=2\delta^{ij}I$ ($i,j=1,2,3$,$I$:4×$\times 4$ 4 unit matrix), $\alpha^{i}\beta+\beta\alpha^{i}=0$ and $\beta^{2}=I$.
@@
The following quantization conditions are imposed on field operators with spin 1/2,
@@
Here, for the sake of completeness, we comment on a boson with spin 0. For a free complex scalar particle $\phi$ with a mass $m$, the Lagrangian density in QFT is given by
^$\widehat{\mathscr{L}}_{\phi}=\hbar c\left\{\partial_{\mu}\widehat{\phi}^{\dagger}(\bm{x},t)\partial^{\mu}\widehat{\phi}(\bm{x},t)-\left(\frac{mc}{\hbar}\right)^{2}\widehat{\phi}^{\dagger}(\bm{x},t)\widehat{\phi}(\bm{x},t)\right\},$ (18)
where $\widehat{\phi}(\bm{x},t)$ is the quantum field of $\phi$. The canonical conjugate of $\widehat{\phi}(\bm{x},t)$ and $\widehat{\phi}^{\dagger}(\bm{x},t)$ are defined by
$\widehat{\pi}_{\phi}(\bm{x},t)\equiv\frac{\partial\widehat{\mathscr{L}}_{\phi}}{\partial(\partial\widehat{\phi}(\bm{x},t)/\partial t)}=\frac{\hbar}{c}\frac{\partial}{\partial t}\widehat{\phi}^{\dagger}(\bm{x},t),~{}~{}\widehat{\pi}^{\dagger}_{\phi}(\bm{x},t)\equiv\frac{\partial\widehat{\mathscr{L}}_{\phi}}{\partial(\partial\widehat{\phi}^{\dagger}(\bm{x},t)/\partial t)}=\frac{\hbar}{c}\frac{\partial}{\partial t}\widehat{\phi}(\bm{x},t),$ (19)
respectively, and the Hamiltonian operator $\widehat{H}_{\phi}$ is obtained as
$\widehat{H}_{\phi}=\int\hbar c\left\{\frac{1}{\hbar^{2}}\widehat{\pi}^{\dagger}_{\phi}(\bm{x},t)\widehat{\pi}_{\phi}(\bm{x},t)+\bm{\nabla}\widehat{\phi}^{\dagger}(\bm{x},t)\cdot\bm{\nabla}\widehat{\phi}(\bm{x},t)+\left(\frac{mc}{\hbar}\right)^{2}\widehat{\phi}^{\dagger}(\bm{x},t)\widehat{\phi}(\bm{x},t)\right\}d^{3}x.$ (20)
The following commutation relations are imposed on the field operators with spin 0:
$\left[\widehat{\phi}(\bm{x},t),\widehat{\pi}_{\phi}(\bm{y},t)\right]=i\hbar\delta^{3}(\bm{x}-\bm{y}),~{}~{}\left[\widehat{\phi}(\bm{x},t),\widehat{\phi}(\bm{y},t)\right]=0,~{}~{}\left[\widehat{\pi}_{\phi}(\bm{x},t),\widehat{\pi}_{\phi}(\bm{y},t)\right]=0,$ (21)
$\left[\widehat{\phi}^{\dagger}(\bm{x},t),\widehat{\pi}^{\dagger}_{\phi}(\bm{y},t)\right]=i\hbar\delta^{3}(\bm{x}-\bm{y}),~{}~{}\left[\widehat{\phi}^{\dagger}(\bm{x},t),\widehat{\phi}^{\dagger}(\bm{y},t)\right]=0,~{}~{}\left[\widehat{\pi}^{\dagger}_{\phi}(\bm{x},t),\widehat{\pi}^{\dagger}_{\phi}(\bm{y},t)\right]=0,$
$\left[\widehat{\phi}(\bm{x},t),\widehat{\pi}^{\dagger}_{\phi}(\bm{y},t)\right]=0,~{}~{}\left[\widehat{\phi}^{\dagger}(\bm{x},t),\widehat{\pi}_{\phi}(\bm{y},t)\right]=0,~{}~{}\left[\widehat{\phi}(\bm{x},t),\widehat{\phi}^{\dagger}(\bm{y},t)\right]=0,$
$\left[\widehat{\pi}_{\phi}(\bm{x},t),\widehat{\pi}^{\dagger}_{\phi}(\bm{y},t)\right]=0.$
In the Schrödinger picture, quantum fields are independent of time, and $\widehat{\pi}_{\phi}(\bm{y})$ and $\widehat{\pi}^{\dagger}_{\phi}(\bm{y})$ are represented by the functional derivatives such as $\widehat{\pi}_{\phi}(\bm{y})=-i\hbar{\delta}/{\delta{\phi}(\bm{y})}$ and $\widehat{\pi}^{\dagger}_{\phi}(\bm{y})=-i\hbar{\delta}/{\delta{\phi}^{\dagger}(\bm{y})}$, respectively. A wave functional $\varPsi(\phi,t)=\langle\phi|\varPsi(t)\rangle$ is evolved by the Schrödinger equation:
$i\hbar\frac{\partial}{\partial t}\varPsi(\phi,t)=\widehat{H}_{\phi}\varPsi(\phi,t),$ (22)
where the Hamiltonian operator is rewritten as
$\widehat{H}_{\phi}=\int\hbar c\left\{-\frac{\delta^{2}}{\delta{\phi}^{\dagger}(\bm{x})\delta{\phi}(\bm{x})}+\bm{\nabla}{\phi}^{\dagger}(\bm{x})\cdot\bm{\nabla}{\phi}(\bm{x})+\left(\frac{mc}{\hbar}\right)^{2}{\phi}^{\dagger}(\bm{x}){\phi}(\bm{x})\right\}d^{3}x.$ (23)
In a similar way as a particle with spin 1/2 and a particle with spin 0, a particle with spin 1 (gauge boson) and a particle with a higher spin are also described by the use of QFT. Actually, particle physics at the electroweak scale is excellently controlled by the Lagrangian density ^$\widehat{\mathscr{L}}_{\rm SM}$ of the standard model, and the basic formula in the Schrödinger picture is given by
$i\hbar\frac{\partial}{\partial t}\varPsi(\varphi_{\rm SM},t)=\widehat{H}_{\rm SM}\varPsi(\varphi_{\rm SM},t),$ (24)
where $\widehat{H}_{\rm SM}$ is the Hamiltonian operator derived from ^$\widehat{\mathscr{L}}_{\rm SM}$, and $\varphi_{\rm SM}$ denotes a set of field variables of the standard model particles.
Let us come back to the scalar particle $\phi$. When we deal with a relativistic scalar particle in QM, we encounter the problem that the probability interpretation breaks down in the absence of a positive definite expression of probability density. Because QM shows excellence in a description of a system with a definite number of particles, we do well to consider a non-relativistic case. Using the redefinition of quantum field:
$\widehat{\phi}(\bm{x},t)=\sqrt{\frac{\hbar}{2mc}}e^{-\frac{i}{\hbar}mc^{2}t}~{}\widehat{\psi}(\bm{x},t),$ (25)
^$\widehat{\mathscr{L}}_{\phi}$ is rewritten as
^$\widehat{\mathscr{L}}_{\phi}$ $=\frac{i\hbar}{2}\left(\widehat{\psi}^{\dagger}(\bm{x},t)\frac{\partial}{\partial t}\widehat{\psi}(\bm{x},t)-\frac{\partial}{\partial t}\widehat{\psi}^{\dagger}(\bm{x},t)~{}\widehat{\psi}(\bm{x},t)\right)-\frac{\hbar^{2}}{2m}\bm{\nabla}\widehat{\psi}^{\dagger}(\bm{x},t)\cdot\bm{\nabla}\widehat{\psi}(\bm{x},t)+\cdots$
$=\widehat{\psi}^{\dagger}(\bm{x},t)i\hbar\frac{\partial}{\partial t}\widehat{\psi}(\bm{x},t)-\widehat{\psi}^{\dagger}(\bm{x},t)\left(-\frac{\hbar^{2}}{2m}\bm{\nabla}^{2}\right)\widehat{\psi}(\bm{x},t)+({\rm total~{}~{}derivatives})+\cdots,$ (26)
where the ellipsis contains a second time-derivative term and this term is neglected in the non-relativistic limit, i.e., ${|mc^{2}\widehat{\psi}|\gg|i\hbar\partial\widehat{\psi}}/{\partial t}|$. From eq. (26), we find that the Hamiltonian operator has an embedded structure such as ${\widehat{H}_{\psi}=\int\widehat{\psi}^{\dagger}(\bm{x},t)\widehat{H}\widehat{\psi}(\bm{x},t)d^{3}x}$ with ${\widehat{H}=-\frac{\hbar^{2}}{2m}\bm{\nabla}^{2}}$.
@@ 2.2 Derivation of quantum mechanics
using eq. (14), $\langle 0|e^{\frac{i}{\hbar}\widehat{H}_{\varphi}t}=\langle 0|$, i.e., $\langle 0|\widehat{H}_{\varphi}=0$ and eq. (10). Here, according to the ordinary procedure, we redefine $\widehat{H}_{\varphi}$ by using the normal ordering. From eq. (33) and the first equation in eq. (8), we find that $\psi(\bm{x},t)$ satisfies the Schrödinger equation in QM:
@@
from the rotation. In this way, it is ascertained that a nested structure exists for $\widehat{{\bm{P}}}$ and $\widehat{\bm{J}}$ other than the Hamiltonian operator $c\widehat{P}_{0}$, as seen from the formulas (42) and (43).
@@
we arrive at the probability interpretation that $\widetilde{\psi}^{\dagger}(\bm{k},t)\widetilde{\psi}(\bm{k},t)$ is the probability density and $\langle\psi(t)|\widehat{\bm{p}}|\psi(t)\rangle$ is the expectation value of momenta for the one-particle system in QM.
@@
In the Schödinger picture, observables in a system described by a Lagrangian density with a first time-derivative term are, in general, given as a form
@@
with a potential $\widehat{V}_{\rm I}(t)$, using the Born approximation in the non-relativistic limit.
We list central features concerning the relationship between QFT and QM, in a system described by a Lagrangian density with a first time-derivative term.
@@ 3 Quantum field’s functional theory
Let us come back to Q1, i.e., “Why does QFT work as an effective theory of elementary particles extremely well?  Why are particles or fields quantized in the first place?”. In concrete, why is a fermion (boson) described by the field operator such as $\widehat{\pi}(\bm{x})=i\hbar{\delta}/{\delta{\varphi}(\bm{x})}$, i.e., $\widehat{\varphi}^{\dagger}(\bm{x})={\delta}/{\delta{\varphi}(\bm{x})}$ ($\widehat{\pi}(\bm{x})=-i\hbar{\delta}/{\delta{\varphi}(\bm{x})}$), using the functional derivative?
If there were a fundamental framework that reaches QFT in a similar way as the derivation of QM from QFT, we can have an answer to the above questions. To explore such a framework, we will draw on the embedded structure in QFT, e.g.,
@@ 3.1 Framework of quantum field’s functional theory
where $\widehat{\varPhi}(\{\varphi\},t)$ is a functional operator, $\{\varphi\}$ and $t$ stand for a field of $\varphi$ and a time, respectively, and $\widehat{H}_{\varphi}$ is the Hamiltonian operator in QFT containing $\varphi(\bm{x})$ and its functional derivatives $\delta/\delta\varphi(\bm{x})$ (see eq. (16)). For instance, $\widehat{H}_{\varphi}$ is given by
@@
We impose the following quantization conditions on functional operators333 If eigenvalues of $\widehat{H}_{\varphi}$ are bounded below, this system can also be quantized using commutation relations although the Lagrangian consists of a first time-derivative term and a (sign-flipping of) Hamiltonian (see eq. (61)). This feature is different from that of the quantization for particles with spin 1/2 in QFT. In fact, when we impose commutation relations on Dirac or Weyl fermion, the system becomes ill-defined with the advent of negative norm states. It stems from the existence of negative-energy states in relativistic QM.,
@@ 3.2 Derivation of quantum field theory
  • Elementary particles create and annihilate. Particles can appear in the vacuum, and a multi-particle state can return to the vacuum state. The vacuum state plays a role as a basis to construct any physical states, e.g., multi-particle states are obtained by multiplying creation operators.
@@
${\varPsi}(\{\varphi\},t)=\psi^{(0)}(t)+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}~{}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N}),$ (84)
@@
$\widehat{\varPhi}(\{\varphi\},t)=e^{-\frac{i}{\hbar}\mathscr{E}_{0}t}\widehat{\varPhi}(\{\varphi\}),~{}~{}\widehat{\varPhi}^{\dagger}(\{\varphi\},t)=\widehat{\varPhi}^{\dagger}(\{\varphi\})e^{\frac{i}{\hbar}\mathscr{E}_{0}t}.$ (88)
@@
Next let us justify that a fermion (boson) field $\varphi(\bm{x})$ is quantized in order to satisfy the anti-commutation relations (the commutation relations). Under an infinitesimal translation $\bm{x}^{\prime}=\bm{x}-\bm{\varepsilon}$, the field $\varphi(\bm{x})$ transforms as $\varphi^{\prime}(\bm{x}^{\prime})=\varphi^{\prime}(\bm{x}-\bm{\varepsilon})=\varphi(\bm{x})$, and then an infinitesimal change of $\varphi(\bm{x})$ is given by
@@
In the same way, the expectation value of energy for the state given by eq. (83) is calculated as
${}^{(1)}\langle\varPsi_{\rm M}(t)|\widehat{H}_{\varPhi}|\varPsi_{\rm M}(t)\rangle^{(1)}=\langle\varPsi(t)|\widehat{H}_{\varPhi}|\varPsi(t)\rangle={\varPsi}^{\dagger}(\{\varphi\},t)\widehat{H}_{\varphi}{\varPsi}(\{\varphi\},t)$
$~{}~{}~{}~{}=\int\mathscr{D}\varphi~{}{\varPsi}^{\dagger}(\varphi,t)\widehat{H}_{\varphi}{\varPsi}(\varphi,t)=\langle\varPsi(t)|\widehat{H}_{\varphi}|\varPsi(t)\rangle.$ (98)